# Fractal Patterns May Unravel the Intelligence in Next-Token Prediction

### v1

@@ Abstract
We study the fractal structure of language, aiming to provide a precise formalism for quantifying properties that may have been previously suspected but not formally shown. We establish that language is: (1) self-similar, exhibiting complexities at all levels of granularity, with no particular characteristic context length, and (2) long-range dependent (LRD), with a Hurst parameter of approximately $\mathrm{H}=0.70\pm 0.09$. Based on these findings, we argue that short-term patterns/dependencies in language, such as in paragraphs, mirror the patterns/dependencies over larger scopes, like entire documents. This may shed some light on how next-token prediction can lead to a comprehension of the structure of text at multiple levels of granularity, from words and clauses to broader contexts and intents. We also demonstrate that fractal parameters improve upon perplexity-based bits-per-byte (BPB) in predicting downstream performance. We hope these findings offer a fresh perspective on language and the mechanisms underlying the success of LLMs.
Language, Fractals, Self-similarity, long-range dependence, Hurst exponent, scaling laws, LLM
@@ 1 Introduction
How does next-token prediction in large language models (LLMs) yield remarkably intelligent behavior? Consider, for instance, the two models: Gemini (Anil et al., 2023a) and GPT4 (OpenAI, 2023). These models have demonstrated extraordinary capabilities beyond just mastering language. Their skills extend to quantitative reasoning, creative content creation, document summarization, and even coding, which has prompted some researchers to ponder if there was more to intelligence than “on-the-fly improvisation” (Bubeck et al., 2023). While understanding the exceptional capabilities of LLMs is complex, particularly given the fuzzy meaning of “intelligent” behavior, a possible insight can be drawn from the study of fractals and self-similarity. We elucidate this connection in this work.
@@
Self-Similarity. Self-similar processes were introduced by Kolmogorov in 1940 (Kolmogorov, 1940). The notion garnered considerable attention during the late 1960s, thanks to the extensive works of Mandelbrot and his peers (Embrechts & Maejima, 2000). Broadly speaking, an object is called “self-similar” if it is invariant across scales, meaning its statistical or geometric properties stay consistent irrespective of the magnification applied to it (see Figure 1). Nature and geometry furnish us with many such patterns, such as coastlines, snowflakes, the Cantor set and the Kuch curve. Despite the distinction, self-similarity is often discussed in the context of “fractals,” another term popularized by Mandelbrot in his seminal book The Fractal Geometry of Nature (Mandelbrot, 1982). However, the two concepts are different (Gneiting & Schlather, 2004). See Section 2.
@@
In language, in particular, there have been studies arguing for the presence of a self-similar structure. Nevertheless, due to the computational constraints of the past, it was not feasible to holistically model the joint probability distribution of language. As such, linguists often resorted to rudimentary approximations in their arguments, such as by substituting a word with its frequency or length (Ausloos, 2012), or by focusing on the recurrence of a specific, predetermined word (Najafi & Darooneh, 2015; Altmann et al., 2012). These studies fall short of fully capturing the underlying structure of language due to the simplifying assumptions they make, as discussed in Section 4.
@@
Highlighting the self-similar nature of a process can have profound implications. For instance, conventional Poisson models for Ethernet traffic were shown to fail because traffic was self-similar (Crovella & Bestavros, 1995; Leland et al., 1994; Paxson & Floyd, 1995; Willinger et al., 1997). In such cases, recognizing and quantifying this self-similarity had practical applications, such as in the design of buffers in network devices (Leland & Wilson, 1991). Similarly in language, we argue that self-similarity may offer a fresh perspective on the mechanisms underlying the success of LLMs. Consider the illustrative example shown in Figure 1, where the task is to predict the subsequent measurement in a time series, specifically predicting next tokens in a Wikipedia article (see Section 2 for details). The three plots in Figure 1 (top) represent different manifestations of the same process observed across three distinct time scales. Notably, we observe rich details, e.g. burstiness, in all of them. Hence, for the model to successfully predict the next measurement, it must capture the behavior of the process at various levels of granularity. The standard approach for quantifying self-similarity is the Hölder exponent (Watkins, 2019), which we denote by $\mathrm{S}$. In language, we estimate it to be $\mathrm{S}=0.59\pm 0.08$, confirming statistical self-similarity.
Refer to captionRefer to captionRefer to caption     Refer to captionRefer to captionRefer to caption
@@
Figure 1: Manifestations of processes across different time scales. A region marked in red corresponds to the magnified plot above it. top: The process exhibits self-similarity with rich details at all levels of granularity. It is an integral process $(X_{t})_{t\in\mathbb{N}}$ calculated from Wikipedia (see Section 2). bottom: Example of a process that is not self-similar, looking smoother at larger time scales.
@@
Why is this significant? We hypothesize that since LLMs are trained to predict the future of a self-similar process, they develop proficiency in capturing behavior across multiple levels of granularity for two interconnected reasons. First, self-similarity implies that the patterns in language at the level of a paragraph are reflective of the patterns seen at the level of a whole text. Hence, recognizing short-term patterns may also aide in learning broader contexts. Second, because language displays detailed, intricate patterns at every level of granularity, it would not be enough to rely only on the immediate context of a sentence to predict the next token. Instead, the model would need to identify patterns at higher levels of granularity; i.e. understand the direction of the argument, and the broader context and intent. It must balance between short- and long-term contexts.  Willinger et al. (1995) and Altmann et al. (2012) argue for self-similarity in language precisely because of this hierarchical nature.
@@
Long-range dependence. However, self-similarity by itself is not sufficient for a predictive model to exhibit anything resembling “intelligent” behavior. In fact, some self-similar processes, despite their intricacy across all levels of granularity, remain entirely unpredictable. A quintessential example is the simple Brownian motion, which is a Wiener process with independent increments. Its discrete analog $B_{n}$ is defined by $B_{n}=\sum_{i=1}^{n}\varepsilon_{i}$, where $\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$. Despite possessing rich details at all granularities, a model trained to predict $B_{n}$ cannot obviously acquire any intelligence since the process itself has independent increments.
@@
Thus, for intelligent behavior to manifest, the process must have some degree of predictability or dependence as well. One classical metric for quantifying predictability in a stochastic process is the Hurst parameter (Hurst, 1951), developed by the hydrologist H. E. Hurst in 1951 while studying the Nile river flooding. It is generally considered to be a robust metric (Willinger et al., 1995), unlike for instance the wavelet estimator (Abry et al., 1995) and the periodogram method (Geweke & Porter-Hudak, 1983) that can be sensitive to errors (Pilgrim & Taylor, 2018). As discussed in Section 2.3, we estimate the Hurst parameter in language to be $\mathrm{H}=0.70\pm 0.09$. For context, $\mathrm{H}$ can only take values in $[0,1]$. A higher value suggests more predictability or persistence in the data, while a lower Hurst parameter indicates more randomness ($\mathrm{H}=0.5$ for completely random systems).
@@
While it is compelling that our estimate of $\mathrm{H}$ in language lies nearly midway between predictability ($\mathrm{H}=1$) and noise ($\mathrm{H}=0.5$), a Hurst parameter of about $0.75$ turns out to occur commonly in nature, including in river discharges, Ethernet traffic, temperatures, precipitation, and tree rings (Crovella & Bestavros, 1995; Feller, 1951; Aref, 1998). For agents that learn from data, such as LLMs, this value is also reminiscent of processing-based theories of curiosity, which suggest that a sweet spot of complexity exists (not too simple, nor too unpredictable) that facilities or accelerates learning (Kidd & Hayden, 2015).
@@
Importantly, predictability and self-similarity together imply long-range dependence (LRD). This follows from the definition of self-similarity, where the patterns at small scales mirror those at larger scales so, for example, the correlations established at micro levels are also pertinent at macro levels. LRD is arguably necessary for intelligence to emerge in a predictive model because processes with only short-range dependence could be forecasted (somewhat trivially) with lookup tables that provide the likelihood of transitions over brief sequences. By contrast, this is not possible in LRD processes whose contexts extend indefinitely into the past.
@@
From a practical standpoint, we demonstrate in Section 3 that incorporating fractal parameters can markedly enhance the ability to predict downstream performance in LLMs, compared to perplexity-based metrics alone, such as bits-per-byte (BPB). Specifically, we introduce a new metric averaging $\mathrm{H}$ with $1/\mathrm{BPB}$ and show that using it to predict downstream performance can increase the adjusted $R^{2}$ from approximately $0.65$ when using solely BPB, to over $0.86$ with the new metric. We do not observe improvements when predicting rankings, however.
@@
    highlight how the fractal structure of language can offer a unique perspective on the intelligent behavior exhibited by LLMs, and provide a precise formalism to quantify properties, such as long-range dependence.
@@
    carry out a comparative study across different model architectures and scales, and different domains, such as ArXiv, GitHub, and Wikipedia, among others.
@@
    connect fractal patterns with learning. Notably, we show that a “median” Hurst exponent (defined in Section 3) improves upon perplexity-based bits-per-byte (BPB) in predicting downstream performance.
@@ 2 Fractal Structure of Language
Figure 2: Peak probability $p_{\epsilon}(\tau)$ is plotted against the granularity level $\tau$ (see Section 2.2). We observe a power law $p_{\epsilon}(\tau)\sim\tau^{-\mathrm{S}}$ in all domains, indicating a self-similar structure, with a median self-similarity exponent of $\mathrm{S}=0.59\pm 0.08$.
@@ 2.1 Preliminaries
Suppose we have a discrete-time, stationary stochastic process $(x_{t})_{t\in\mathbb{N}}$, with $\mathbb{E}[x_{t}]=0$ and $\mathbb{E}[x_{t}^{2}]=1$. We will refer to $(x_{t})_{t\in\mathbb{N}}$ as the increment process to distinguish it from the integral process $(X_{t})_{t\in\mathbb{N}}$ defined by $X_{t}=\sum_{k=0}^{t}x_{k}$. While $(x_{t})_{t\in\mathbb{N}}$ and $(X_{t})_{t\in\mathbb{N}}$ are merely different representations of the same data, it is useful to keep both representations in mind. For example, self-similarity is typically studied in the context of integral processes whereas long-range dependence (LRD) is defined on increment processes.
In the literature, it is not uncommon to mistakenly equate parameters that are generally different. For example, the Hurst parameter has had many different definitions in the past that were not equivalent, and Mandelbrot himself had cautioned against this (Mandelbrot, 2002). The reason behind this is because different parameters can agree in the idealized fractional Brownian motion setting, leading some researchers to equate them in general (Watkins, 2019). We will keep the self-similarity exponent $\mathrm{S}$ and the Hurst parameter $\mathrm{H}$ separate in our discussion.
@@
In order to establish self-similarity and LRD in language, we convert texts into sequences of bits using a large language model (LLM). Specifically, we use PaLM2-L (Unicorn) (Anil et al., 2023b) to calculate the probability of the next token $w_{t}$ conditioned on its entire prefix $w_{[t-1]}=(w_{0},w_{1},\ldots,w_{t-1})$. By the chain rule (Cover, 1999), the corresponding number of bits assigned to $w_{t}$ is $z_{t}=-\log p(w_{t}|w_{[t-1]})$. Unlike in prior works, which rely on simplifications such as by substituting a word with its length (Ausloos, 2012) or by focusing on the recurrence of a single word (Najafi & Darooneh, 2015; Altmann et al., 2012), we use the LLM to approximate the full joint distribution of language. We carry out these calculations for prefixes of up to 2048 tokens ($\approx 8$ pages of text). Since language is a stochastic process, the sequence of bits of each token conditioned on its past converges asymptotically to the average number of bits required to encode the entire sequence (Cover, 1999). Hence, a suitable normalization, such bits-per-byte (BPB), results in a standardized description of text, consistent across tokenizers. BPB is a widely used as a tokenizer-agnostic metric to compare language modeling performance, e.g. for The Pile (Gao et al., 2020).
@@
Besides PaLM2, we also experiment and report on various model sizes of PaLM (Chowdhery et al., 2022) and decoder-only T5 (Raffel et al., 2019). Namely, we report results for models: PaLM2 XXS (Gecko), XS (Otter), S (Bison), M, and L (Unicorn); PaLM 8B, 62B, 540B; and decoder-only T5.1.1 at Base (110M), Large (341M), XL (1.2B), and XXL (5B) sizes. For PaLM and PaLM2, we use the checkpoints pretrained in Chowdhery et al. (2022) and Anil et al. (2023b). All T5.1.1 decoder baselines, on the other hand, are trained with a casual language modeling objective for 262B tokens of C4 (Raffel et al., 2019). More details on how we train our T5.1.1 baselines can be found in Appendix A.
To rely on LLM for such analysis, it must provide probability scores that are reasonably well-calibrated. Generally, LLMs are known to produce calibrated probability scores at the token level (Kadavath et al., 2022). In Figure 3, we reconfirm this by comparing the logits, $-\log p(\mathrm{word})$, predicted by one of the small language models we use in our study (PaLM-8B) with the actual log probabilities derived from the Google Web Trillion Word Corpus (Brants & Franz, 2006) based on word frequencies. We use histogram binning (by grouping similar logits together) and plot their averaged actual log probabilities, similar to how the expected calibration error (ECE) is calculated (Guo et al., 2017). Notably, we find a strong agreement for the most frequently occurring words, i.e., when the word probability exceeds $p\gg 10^{-9}$.
Figure 3: Comparison of PaLM-8B’s logits with actual log-probabilities. We observe a substantial agreement except for exceedingly uncommon words with a probability $<10^{-9}$. This is consistent with reported findings that LLMs produce calibrated probability scores for tokens; e.g. (Kadavath et al., 2022).
@@
Once $z_{t}$ is computed for a document, we construct the increment process $(x_{t})_{t\in\mathbb{N}}$ by normalizing $z_{t}$ to have a zero-mean and unit variance. The integral process $(X_{t})_{t\in\mathbb{N}}$ is calculated based on $(x_{t})_{t\in\mathbb{N}}$, as described earlier and depicted in Figure 1 (top). Normalizing bits (to have zero mean and unit variance) models language as a random walk. It is a standard approach used extensively in the literature in various contexts, such as in DNA sequences (Peng et al., 1992; Roche et al., 2003; Montemurro & Pury, 2002; Kokol & Podgorelec, 2000; Schenkel et al., 1993).
@@
For analysis, we use The Pile validation split (Gao et al., 2020), consisting of 22 subdomains such as Wikipedia and GitHub. We restrict analysis to sufficiently-long documents of length $>4K$ tokens and use the first 2K tokens only, to sidestep potential effects of the finite length of documents and the model context. To mitigate noise, only domains with $>1K$ documents are compared; we report results for them separately and their median. We use bootstrapping (Efron & Tibshirani, 1994) to estimate the error margin.
@@
Notation. We write $f(x)\sim x^{c}$ if $f(x)=x^{c}L(x)$ for some slowly-varying function $L$; i.e. $L(tx)/L(x)\to 1$ as $x\to\infty$ for all $t>0$. Examples of slowly varying functions are constants $L(x)=c$ and $L(x)=\log x$. When $f(x)\sim x^{c}$, we will abuse terminology slightly by referring to $f(x)$ as a power law function.

### v2

# Fractal Patterns May Illuminate the Success of Next-Token Prediction

@@ Abstract
We study the fractal structure of language, aiming to provide a precise formalism for quantifying properties that may have been previously suspected but not formally shown. We establish that language is: (1) self-similar, exhibiting complexities at all levels of granularity, with no particular characteristic context length, and (2) long-range dependent (LRD), with a Hurst parameter of approximately $\mathrm{H}=0.70\pm 0.09$. Based on these findings, we argue that short-term patterns/dependencies in language, such as in paragraphs, mirror the patterns/dependencies over larger scopes, like entire documents. This may shed some light on how next-token prediction can capture the structure of text across multiple levels of granularity, from words and clauses to broader contexts and intents. In addition, we carry out an extensive analysis across different domains and architectures, showing that fractal parameters are robust. Finally, we demonstrate that the tiny variations in fractal parameters seen across LLMs improve upon perplexity-based bits-per-byte (BPB) in predicting their downstream performance. We hope these findings offer a fresh perspective on language and the mechanisms underlying the success of LLMs.
@@ 1 Introduction
How does the training objective of predicting the next token in large language models (LLMs) yield remarkable capabilities? Consider, for instance, the two models: Gemini [5] and GPT4 [49]. These models have demonstrated capabilities that extend to quantitative reasoning, summarization, and even coding, which has led some researchers to ponder if there was more to intelligence than “on-the-fly improvisation” [11]. While providing a satisfactory explanation is a difficult endeavor, a possible insight can be drawn from fractals and self-similarity. We elucidate the connection in this work.
@@
Self-Similarity. Self-similar processes were introduced by Kolmogorov in 1940 [36]. The notion garnered considerable attention during the late 1960s, thanks to the extensive works of Mandelbrot and his peers [19]. Broadly speaking, an object is called “self-similar” if it is invariant across scales, meaning its statistical or geometric properties stay consistent irrespective of the magnification applied to it (see Figure 1). Nature and geometry furnish us with many such patterns, such as coastlines, snowflakes, the Cantor set and the Kuch curve. Despite the distinction, self-similarity is often discussed in the context of “fractals,” another term popularized by Mandelbrot in his seminal book The Fractal Geometry of Nature [44]. However, the two concepts are different [26]. See Section 2.
@@
In language, in particular, there have been studies arguing for the presence of a self-similar structure. Nevertheless, due to computational constraints, it was not feasible to holistically model the joint probability distribution of language. As such, linguists often resorted to rudimentary approximations in their arguments, such as by substituting a word with its frequency or length [9], or by focusing on the recurrence of a specific, predetermined word [48, 3]. These studies fall short of fully capturing the structure of language due to the simplifying assumptions they make, as discussed in Section 4.
@@
Highlighting the self-similar nature of a process can have profound implications. For instance, conventional Poisson models for Ethernet traffic were shown to fail because traffic was self-similar [16, 38, 50, 66]. In such cases, recognizing and quantifying self-similarity had practical applications, such as in the design of buffers [39]. Similarly in language, we argue that self-similarity may offer a fresh perspective on the mechanisms underlying the success of LLMs. Consider the illustrative example shown in Figure 1, where the task is to predict the subsequent measurement in a time series, specifically predicting next tokens in a Wikipedia article (see Section 2 for details). The three plots in Figure 1 (left) represent different manifestations of the same process observed across three distinct time scales. Notably, we observe rich, self-similar details, such as burstiness, in all of them. A well-established approach for quantifying self-similarity is the Hölder exponent [64], which we denote by $\mathrm{S}$. In language, we find it to be $\mathrm{S}=0.59\pm 0.08$, confirming statistical self-similarity.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
@@
Figure 1: Manifestations of processes across different time scales. A region marked in red corresponds to the magnified plot beneath it. left: The process exhibits self-similarity with rich details at all levels of granularity. It is an integral process $(X_{t})_{t\in\mathbb{N}}$ calculated from Wikipedia (see Section 2). right: Example of a process that is not self-similar, looking smoother at larger time scales.
@@
Why is this important? We hypothesize that since LLMs are trained to predict the future of a self-similar process, they develop proficiency in capturing patterns across multiple levels of granularity for two interconnected reasons. First, self-similarity implies that the patterns at the level of a paragraph are reflective of the patterns seen at the level of a whole text, which is reminiscent of the recursive structure of language [52]. Thus, recognizing short-term patterns can aide in learning broader contexts. Second, because language displays intricate patterns at all levels of granularity, it would not be enough to rely only on the immediate context of a sentence to predict the next token. Instead, the model needs to identify patterns at higher levels of granularity; e.g. follow the direction of the argument and the broader intent. It must balance between short- and long-term contexts.  Willinger et al., [65] and Altmann et al., [3] argue for self-similarity in language due to this hierarchical nature.
@@
Long-range dependence. However, self-similarity alone is not sufficient for a predictive model to exhibit anything resembling “intelligent” behavior. In fact, some self-similar processes, despite their intricate details, remain entirely unpredictable. A quintessential example is the simple Brownian motion, which is a Wiener process with independent increments. Its discrete analog is $B_{n}=\sum_{i=1}^{n}\varepsilon_{i}$, where $\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$. Despite possessing rich details at all granularities, a model trained to predict $B_{n}$ cannot learn anything useful from data since the process itself has independent increments.
@@
Thus, for strong capabilities to emerge, the process must have some degree of predictability or dependence as well. One classical metric for quantifying predictability in a stochastic process is the Hurst parameter [31], developed by the hydrologist H. E. Hurst in 1951 while studying the Nile river. It is generally considered to be a robust metric [65], unlike the wavelet estimator [1] and the periodogram method [24] that can be sensitive to errors [53]. As discussed in Section 2.3, we find the Hurst parameter in language to be $\mathrm{H}=0.70\pm 0.09$. For context, $\mathrm{H}$ only takes values in $[0,1]$. A value $\mathrm{H}>0.5$ implies predictability in the data, while $\mathrm{H}=0.5$ indicates random increments.
@@
While it is compelling that our estimate of $\mathrm{H}$ in language lies nearly midway between predictability ($\mathrm{H}=1$) and noise ($\mathrm{H}=0.5$), a Hurst parameter of about $0.75$ turns out to occur commonly in nature, including in river discharges, Ethernet traffic, temperatures, precipitation, and tree rings [16, 21, 8]. For agents that learn from data, such as LLMs, this value is also reminiscent of processing-based theories of curiosity, which suggest that a sweet spot of complexity exists (not too simple, nor too unpredictable) that facilities or accelerates learning [34].
@@
Importantly, predictability and self-similarity together imply long-range dependence (LRD). This follows from the definition of self-similarity, where the patterns at small scales mirror those at larger scales so, for example, the correlations established at micro levels are also pertinent at macro levels. LRD is arguably crucial for enhancing the functionality of predictive models because processes with only short-range dependence could be forecasted (somewhat trivially) with lookup tables that provide the likelihood of transitions over brief sequences. By contrast, this is not possible in LRD processes whose contexts extend indefinitely into the past.
Information Theoretic Complexity.
To define fractal parameters, we follow recent works such as [28, 22, 40, 46, 25] in adopting an information-theoretic characterization of the complexity in language using minimal-length codes or surprise. This corresponds to an intrinsic, irreducible description of language and the minimum compute overhead to comprehend/decode it [22], which also correlates well with actual reading times [28, 40]. In this context, self-similarity means that the intrinsic complexity or surprise in language (measured in bits) cannot be smoothed out, even as we look into broader narratives. That is, surprising paragraphs will follow predictable paragraphs, in a manner that is statistically similar to how surprising sentences follow predictable sentences.
Analysis.
How robust are these findings? To answer this question, we carry out an extensive empirical analysis across various model architectures and scales, ranging from 1B to over 500B parameters. We find that fractal parameters are quite robust to the choice of the architecture.
@@
However, there exists tiny variations across LLMs. Interestingly, we demonstrate that from a practical standpoint, these differences help in predicting downstream performance in LLMs compared to using perplexity-based metrics alone, such as bits-per-byte (BPB). Specifically, we introduce a new metric and show that using it to predict downstream performance can increase the adjusted $R^{2}$ from approximately $0.65$ when using solely BPB, to over $0.86$ with the new metric [^1] We release the code for calculating fractal parameters at: https://github.com/google-research/google-research/tree/master/fractals_language .
@@
    highlight how the fractal structure of language can offer a new perspective on the capabilities of LLMs, and provide a formalism to quantify properties, such as long-range dependence.
@@
    carry out a comparative study across different model architectures and scales, and different domains, such as ArXiv and GitHub, demonstrating that fractal parameters are robust.
@@
    connect fractal patterns with learning. Notably, we show that a “median” Hurst exponent improves upon perplexity-based bits-per-byte (BPB) in predicting downstream performance.
@@ 2.1 Preliminaries
Suppose we have a discrete-time, stationary stochastic process $(x_{t})_{t\in\mathbb{N}}$, with $\mathbb{E}[x_{t}]=0$ and $\mathbb{E}[x_{t}^{2}]=1$. We will refer to $(x_{t})_{t\in\mathbb{N}}$ as the increment process to distinguish it from the integral process $(X_{t})_{t\in\mathbb{N}}$ defined by $X_{t}=\sum_{k=0}^{t}x_{k}$. While $(x_{t})_{t\in\mathbb{N}}$ and $(X_{t})_{t\in\mathbb{N}}$ are merely different representations of the same data, it is useful to keep both representations in mind. For example, self-similarity is typically studied in the context of integral processes whereas LRD is defined on increment processes.
In the literature, it is not uncommon to mistakenly equate parameters that are generally different. For example, the Hurst parameter $\mathrm{H}$ has had many definitions in the past that were not equivalent, and Mandelbrot himself cautioned against this [43]. The reason behind this is because different parameters can agree in the idealized fractional Brownian motion, leading some researchers to equate them in general [64]. We will keep the self-similarity exponent $\mathrm{S}$ and $\mathrm{H}$ separate in our discussion.
Figure 2: Peak probability $p_{\epsilon}(\tau)$ is plotted against the granularity level $\tau$ (see Section 2.2). We observe power laws $p_{\epsilon}(\tau)\sim\tau^{-\mathrm{S}}$, indicating self-similarity, with a median exponent of $\mathrm{S}=0.59\pm 0.08$.
@@
In order to establish self-similarity and LRD in language, we convert texts into sequences of bits using a large language model (LLM). Specifically, we use PaLM2-L (Unicorn) [6] to calculate the probability of the next token $w_{t}$ conditioned on its entire prefix $w_{[t-1]}=(w_{0},w_{1},\ldots,w_{t-1})$. As discussed in Section 1, this captures its intrinsic, irreducible description [22]. By the chain rule [15], the corresponding number of bits assigned to $w_{t}$ is $z_{t}=-\log p(w_{t}|w_{[t-1]})$. Unlike in prior works, which rely on simplifications such as by substituting a word with its length [9] or by focusing on the recurrence of a single word [48, 3], we use the LLM to approximate the full joint distribution of language since LLMs are known to produce calibrated probability scores at the token level [33]. We carry out these calculations for prefixes of up to 2048 tokens ($\approx 8$ pages of text). With a suitable normalization, such bits-per-byte (BPB), one obtains a standardized description of text, consistent across tokenizers. BPB is a widely used as a tokenizer-agnostic metric to compare LM modeling performance, e.g. for The Pile [23].
Figure 3: Rescaled range $R(n)/S(n)$ is plotted against the number of normalized bits $n$. We observe a power law $R(n)/S(n)\sim n^{\mathrm{H}}$ in all domains. When aggregating all datasets, $\mathrm{H}=0.70\pm 0.09$.
@@
Besides PaLM2, we also experiment and report on various model sizes of PaLM [12] and decoder-only T5 [54]. Namely, we report results for models: PaLM2 XXS (Gecko), XS (Otter), S (Bison), M, and L (Unicorn); PaLM 8B, 62B, 540B; and decoder-only T5.1.1 at Base (110M), Large (341M), XL (1.2B), and XXL (5B) sizes. For PaLM and PaLM2, we use the checkpoints pretrained in Chowdhery et al., [12] and Anil et al., 2023b [6]. All T5.1.1 decoder baselines, on the other hand, are trained with a casual language modeling objective for 262B tokens of C4 [54]. All experiments are executed on Tensor Processing Units (TPUs). More details on how we train T5.1.1 baselines are in Appendix A.
@@
Once $z_{t}$ is computed for a document, we follow standard definitions in constructing the increment process $(x_{t})_{t\in\mathbb{N}}$ by normalizing $z_{t}$ to have a zero-mean and unit variance. Intuitively, fractal parameters are intended to measure a fundamental property of the process (e.g. LRD) that should not be affected by scale, hence the normalization. The integral process $(X_{t})_{t\in\mathbb{N}}$ is calculated based on $(x_{t})_{t\in\mathbb{N}}$, as described earlier and depicted in Figure 1 (top). Normalizing bits (to have zero mean and unit variance) models language as a random walk. It is a standard approach used extensively in the literature in various contexts, such as in DNA sequences [51, 56, 47, 35, 58].
@@
For analysis, we use The Pile validation split [23], consisting of 22 subdomains such as Wikipedia and GitHub. We restrict analysis to sufficiently-long documents of length $>4K$ tokens and use the first 2K tokens only, to sidestep potential effects of the finite length of documents and the model context. To mitigate noise, only domains with $>1K$ documents are compared; we report results for them separately and their median. We use bootstrapping [17] to estimate the error margin.
@@
Notation. We write $f(x)\sim x^{c}$ if $f(x)=x^{c}L(x)$ for some function $L$ that satisfies $L(tx)/L(x)\to 1$ as $x\to\infty$ for all $t>0$. Examples of slowly varying functions are constants $L(x)=c$ and $L(x)=\log x$. When $f(x)\sim x^{c}$, we abuse terminology slightly by referring to $f(x)$ as a power law.
Refer to captionRefer to caption
Figure 4: left: Estimates of the self-similarity exponent $\mathrm{S}$ are generally robust to the choice of $\epsilon$. right: The partial auto-correlation function calculated across domains. DM Mathematics has a much shorter dependence compared to the rest of the domains, in agreement with its Hurst parameter.